OpenStudy (anonymous):

Metal award :33 A car travels a distance of 70 miles in 2, 1/2 hrs. How much faster on the average must it travel to make the trip in 3/4 hrs less time? (MPH)

OpenStudy (anonymous):

70 miles/2.5 hrs= 28 mph 3/4 hours less= 2.5-0.75= 1.75 hrs 70 miles/ 1.75= 40 mph This is a difference of 40-28= 12 mph.

OpenStudy (anonymous):

Thank you so much!!! ^^